You are planning to accumulate ¥14,000,000 to purchase a home in Tokyo in 8 years with...

  1. You are planning to accumulate ¥14,000,000 to purchase a home in Tokyo in 8 years with Dia-Ichi Kangyo Bank. For the next 8 years, you earn an annual return of 12% compounded semiannually. How much should you deposit with Dia-Ichi Kangyo Bank every six months?
Answer #1

Future value = pmtx ((1+r)^n-1))/ PMT = FV = ((1+r)^n-1)/ Save each six months = PMT = 545,330.01 [ 14,000,000/(((1.06-16)-1)